# IPAD: Inverse Prompt for AI Detection – A Reliable and Explainable LLM-Generated Essay Detector

### v2

@@ Abstract
Large Language Models (LLMs) have attained human-level fluency in text generation, which complicates the distinction between human-written and LLM-generated texts. This increases the risk of misuse and highlights the need for reliable detectors. Yet, existing detectors exhibit poor robustness on out-of-distribution (OOD) data and attacked data, which is critical for real-world scenarios. Also, they struggle to provide interpretable evidence to support their decisions, thus undermining the reliability. In light of these challenges, we propose  IPAD (Inverse Prompt for AI Detection), a novel framework consisting of a  Prompt Inverter that identifies predicted prompts that could have generated the input text, and two  Distinguishers that examine the probability that the input texts align with the predicted prompts. Empirical evaluations demonstrate that IPAD outperforms the strongest baselines by 9.05% (Average Recall) on in-distribution data, 12.93% (AUROC) on out-of-distribution (OOD) data, and 5.48% (AUROC) on attacked data. IPAD also performs robustly on structured datasets. Furthermore, an interpretability assessment is conducted to illustrate that IPAD enhances the AI detection trustworthiness by allowing users to directly examine the decision-making evidence, which provides interpretable support for its state-of-the-art detection results.
IPAD: Inverse Prompt for AI Detection – A Reliable and Explainable LLM-Generated Essay Detector
Zheng Chen1thanks: These authors contributed equally to this work. , Yushi Feng211footnotemark: 1 , Yue Deng1 , Changyang He3 , Hongxi Pu4,Bo Li1 1 Computer Science and Engineering, Hong Kong University of Science and Technology 2 School of Computing and Data Science, The University of Hong Kong 3 Max Planck Institute for Security and Privacy 4 Computer Science, The University of Michigan zchenin@connect.ust.hk, fengys@connect.hku.hk, ydengbi@connect.ust.hk, changyang.he@mpi-sp.org, hongxi@umich.edu, bli@cse.ust.hk
@@ 1 Introduction
Large Language Models (LLMs), characterized by their massive scale and extensive training data  (Chen et al., 2024), have achieved significant advances in natural language processing (NLP)  (Ouyang et al., 2022; Veselovsky et al., 2023; Wu et al., 2025). However, with the advanced capabilities of LLMs, they are subject to frequent misused in various domains, including academic fraud, the creation of deceptive material, and the generation of fabricated information (Ji et al., 2023; Pagnoni et al., 2022; Mirsky et al., 2023), which underscores the critical need to distinguish between human-written text (HWT) and LLM-generated text (LGT) (Pagnoni et al., 2022; Yu et al., 2025; Kirchenbauer et al., 2023).
However, due to their sophisticated functionality, LLMs pose significant challenges in the robustness of current AI detection systems (Wu et al., 2025). The existing detection systems, including commercial ones, frequently misclassify texts as HWT (Price and Sakellarios, 2023; Walters, 2023) and generate inconsistent results when analyzing the same text using different detectors (Chaka, 2023; Weber-Wulff et al., 2023). Studies show false positive rates reaching up to 50% and false negative rates as high as 100% in different tools (Weber-Wulff et al., 2023) when dealing with out-of-distribution (OOD) datasets.
Another critical issue with the existing AI detection systems is their lack of verifiable evidence (Halaweh and Refae, 2024), as these tools typically provide only simple outputs like "likely written by AI" or percentage-based predictions (Weber-Wulff et al., 2023). The lack of evidence prevents users from defending themselves against false accusations (Chaka, 2023) and hinders organizations from making judgments based solely on the detection results without convincing evidences (Weber-Wulff et al., 2023). This problem is particularly troublesome not only because the low accuracy of such systems as mentioned before, but also due to the consequent inadequate response to LLM misuse, which can lead to significant societal harm (Stokel-Walker and Van Noorden, 2023; Porsdam Mann et al., 2023; Shevlane et al., 2023; Wu et al., 2025). These limitations highlight the pressing need for more reliable, explainable and robust detectors.
@@
Empirical results show that IPAD outperforms state-of-the-art baselines by 9.05% in Average Recall on in-distribution datasets, 12.93% in AUROC on out-of-distribution (OOD) datasets, and 5.48% in AUROC under adversarial attacks. IPAD also generalizes well to structured data. A user study further reveals that IPAD improves trust and usability in detection tasks by presenting concrete decision evidence, including predicted prompts and regenerated texts. Code is available [^1] https://github.com/Bellafc/IPAD-Inver-Prompt-for-AI-Detection-/.
@@ 2.1 Preliminaries
Our method comprises a Prompt Inverter  $f_{\text{inv}}$, and two Distinguishers, namely the Prompt-Text Consistency Verifier (PTCV) $f_{\text{PTCV}}$ and the Regeneration Comparator (RC) $f_{\text{RC}}$. Given an input text $T$, the task is to determine whether it is human-written (HWT) or generated by an LLM (LGT). We denote by $\mathcal{D}_{\text{PI}}$ the training set for $f_{\text{inv}}$, consisting of pairs $(T,P)$ where $T$ is an LLM-generated text and $P$ is its original prompt. The two distinguishers are trained using disjoint datasets: $\mathcal{D}_{\text{LGT}}$ contains LLM-generated samples and $\mathcal{D}_{\text{HWT}}$ contains human-written ones. All components are fine-tuned using Microsoft’s Phi3-medium-128k-instruct model. [^2] https://huggingface.co/microsoft/Phi-3-medium-128k-instruct.
@@
To estimate the fine-tuned model’s binary classification probability (i.e., the probability of predicting “yes” or “no”), we follow the logit-based estimation approach (Yoshikawa and Okazaki, 2023). Given the model input $x$, and the output logits $z$, the model’s probability assigned to $\hat{y}$ is computed through the softmax function $\sigma$ :
$\text{Confidence}_{\text{yes}}=P(\hat{y}=\text{``yes''}\mid x)=\sigma(z)_{\text{yes}},\quad$
$\text{Confidence}_{\text{no}}=P(\hat{y}=\text{``no''}\mid x)=\sigma(z)_{\text{no}}$
@@
$\text{Probability}_{\text{yes}}=\frac{\text{Confidence}_{\text{yes}}}{\text{Confidence}_{\text{yes}}+\text{Confidence}_{\text{no}}},\quad$
$\text{Probability}_{\text{no}}=1-\text{Probability}_{\text{yes}}$
@@ 2.2 Workflow
We perform inference on unseen input texts $T$ by sequentially applying the trained modules. Given an input text $T$, we first use the prompt inverter $f_{\text{inv}}$ to recover the most plausible prompt $P$. The prompt is then used to regenerate a candidate text $T^{\prime}$ via the an LLM. Next, we compute two probability scores: $p_{\text{PTCV}}$, indicating whether $T$ is consistent with $P$, and $p_{\text{RC}}$, assessing the likelihood that $T$ and $T^{\prime}$ originate from the same prompt. These scores are fused into a final decision score $\hat{p}$ using the gird-searched weight $w$, and the predicted label is determined by comparing $\hat{p}$ against the threshold $\tau$. The complete algorithm can be found in Appendix  B.
The inference procedure of the IPAD framework consists of three calls through a light-weight open-sourced LLM phi-3-medium-128k-instruct. Phi-3 is a decoder-only Transformer, whithin which, the self-attention complexity per layer is $\mathcal{O}(n^{2}\cdot d)$, where $n$ is the sequence length and $d$ is the hidden dimension (Vaswani et al., 2017). The additional api call to gpt-3.5-turbo for re-generating texts introduces fixed latency but no local computation cost. Therefore, the overall computational cost is bounded by $\mathcal{O}(3\cdot L\cdot n^{2}\cdot d+\texttt{OpenAI}_{\text{api}})$, where $L=32$ is the number of layers in phi-3 (Abdin et al., 2024), which is relatively small. All three phi-3 calls can be deployed in an Nvidia V100 GPU as the minimum requirement. This demonstrates that IPAD is not computationally expensive and can be deployed with relatively modest hardware requirements.
@@ 2.3 Training
The supervised fine-tuning  (Wei et al., 2022) process is performed on a Microsoft’s open model, phi3-medium-128k-instruct, and we use low-rank adaptation (LoRA) method  (Hu et al., 2022) on the LLaMA-Factory framework [^3] https://huggingface.co/papers/2403.13372 (Zheng et al., 2024). We train it using six A800 GPUs for 20 hours for Prompt Inverter, 7 hours for PTCV, and 9 hours for RC.
@@ 3.1.1 Evaluation Baselines and Metrics
The in-distribution experiments refer to the testing results presented in  (Koike et al., 2024), where the data aligns with the training data used for the IPAD, thereby serving as our baseline. This baseline assesses how the RoBERTa classifiers (base and large) (Park et al., 2021), the HC3 detector (Guo et al., 2023), and the OUTFOX detector (Koike et al., 2024) perform on standard data as well as under DIPPER (Alkanhel et al., 2023) and OUTFOX attacks.
The OOD experiments refer to the DetectRL baseline  (Wu et al., 2024), which is a comprehensive benchmark, which includes four datasets: (1) academic abstracts from the arXiv Archive (covering the years 2002 to 2017) [^4] http://kaggle.com/datasets/spsayakpaul/arxiv-paper-abstracts/data, (2) news articles from the XSum dataset  (Narayan et al., 2018), (3) creative stories from Writing Prompts  (Fan et al., 2018), and (4) social reviews from Yelp Reviews (Zhang et al., 2015). It also employs three attack methods to simulate complex real-world detection scenarios, which include (1) the prompt attacks, (2) paraphrase attacks, and (3) perturbation attacks (Wu et al., 2024). DetectRL evaluates three classifiers on the OOD dataset: DetectLLM (LRR) (Su et al., 2023), Fast-DetectGPT (Bao et al., 2023), RoBERTa Classifier (Base). We included two more strong classifiers in our evaluation DetectLLM (NPR) (Su et al., 2023) and Binoculars  (Hans et al., 2024). All the testing sets have 1,000 samples in our experiments.
We further evaluate its performance on OOD datasets with structured prompts. The LongWriter dataset (Bai et al., 2025), featuring an average prompt length of 1,501 tokens, reflects IPAD’s capability to handle long-form prompts. The Code-Feedback [^5] https://huggingface.co/datasets/m-a-p/Code-Feedback and Math datasets (Hendrycks et al., 2021) contain highly structured prompts, in contrast to typical essay-like writing. We compare IPAD with baseline detectors from DetectRL to assess its relative performance under these challenging conditions.
The  Area Under Receiver Operating Characteristic curve (AUROC) is widely used for assessing detection method  (Mitchell et al., 2023). Since our models predict binary labels, we follow the  Wilcoxon-Mann-Whitney statistic (Calders and Jaroszewicz, 2007), and the formula is shown in Appendix  C. The  AvgRec is the average of  HumanRec and  MachineRec, which refers to the recall of the Human-written texts and the LLM-generated texts (Li et al., 2024).
@@ 3.1.3 Comparison of IPAD with other detectors in and out of distribution
Figure 2: The In-distribution data performance of IPAD and the baseline detectors. Since  Koike et al. (2024) only presents the AvgRec data for the baselines, we also calculate AvgRec data for IPAD to compare.
@@
Table 2: Detection Accuracy (AUROC %) on four diverse OOD datasets
@@
Table 3: Detection Accuracy (AUROC %) on three attacked OOD datasets
@@ 3.2.1 Necessity
To prove that it is necessary to fine-tune on IPAD with IPAD with  PTCV and  RC, we conducted ablation study to use the same finetune method on only  input texts and only  predicted prompts, with the finetune data format shown in Appendix  D. We only experimented on Prompt Inverter + PTCV and Prompt Inverter + RC to compare with the three-moduled IPAD.
@@ 3.2.2 Effectiveness
We use DPIC (Yu et al., 2024) and PE (Zhang et al., 2024b) as baseline methods for prompt extraction. DPIC employs a zero-shot approach using the prompt states in Appendix  E, while PE uses adversarial attacks to recover system prompts. In our evaluation, we tested 1000 LGT and 1000 HWT samples. We use only in-distribution data for testing since only these datasets include original prompts. The metrics are all tested on comparing the similarity of the original prompts and the predicted prompts. The results shown in Table  5 illustrate that IPAD consistently outperforms both DPIC and PE across all four metrics (BartScore (Yuan et al., 2021), Sentence-Bert Cosine Similarity (Reimers and Gurevych, 2019), BLEU (Papineni et al., 2002), and ROUGE-1 (Lin, 2004)), which highlight the effectiveness of the IPAD  Prompt Inverter.
@@
We conducted a comparison study using the frozen Prompt Inverter but different distinguishing methods. The first and second methods employed Sentence-Bert (Reimers and Gurevych, 2019) and Bart-large-cnn (Yuan et al., 2021) to compute the similarity score between the input texts and the regenerated texts. We selected thresholds that maximized AvgRec, which were 0.67 for Sentence-Bert and -2.52 for Bart-large-cnn. The classification rule is that the texts with scores greater than the threshold will be classified as LGT, while the texts with scores less than or equal to the threshold will be classified as HWT. The third method is to directly prompt ChatGPT in Appendix  D, which mimic the fine-tuning process of PTCV and RC. The final results shown in Table  6 demonstrate that the other distinguishing methods performed worse than IPAD, highlighting the superior effectiveness of the IPAD  Distinguishers.
@@
DPIC first uses a zero-shot prompt inverter to generate prompts, then applies a Siamese encoder and classifier to measure similarity between the embeddings of the original and regenerated texts. However, the classifier’s reliance on embedding similarity is ambiguous, as similar texts may stem from different prompts. IPAD addresses this by fine-tuning directly on raw texts and reformulating the task as a logical reasoning problem as shown in the instructions of PTCV and RC. Our trained Prompt Inverter outperforms DPIC’s generic zero-shot method as shown in Table 5, and IPAD also achieves better performance than DPIC overall, as results shown in Appendix  F.

### v3

# IPAD: Inverse Prompt for AI Detection - A Robust and Interpretable LLM-Generated Text Detector

@@ Abstract
Large Language Models (LLMs) have attained human-level fluency in text generation, which complicates the distinguishing between human-written and LLM-generated texts. This increases the risk of misuse and highlights the need for reliable detectors. Yet, existing detectors exhibit poor robustness on out-of-distribution (OOD) data and attacked data, which is critical for real-world scenarios. Also, they struggle to provide interpretable evidence to support their decisions, thus undermining the reliability. In light of these challenges, we propose  IPAD (Inverse Prompt for AI Detection), a novel framework consisting of a  Prompt Inverter that identifies predicted prompts that could have generated the input text, and two  Distinguishers that examine the probability that the input texts align with the predicted prompts. Empirical evaluations demonstrate that IPAD outperforms the strongest baselines by 9.05% (Average Recall) on in-distribution data, 12.93% (AUROC) on out-of-distribution data, and 5.48% (AUROC) on attacked data. IPAD also performs robustly on structured datasets. Furthermore, an interpretability assessment is conducted to illustrate that IPAD enhances the AI detection trustworthiness by allowing users to directly examine the decision-making evidence, which provides interpretable support for its state-of-the-art detection results.
@@ 1 Introduction
Large Language Models (LLMs), characterized by their massive scale and extensive training data  (r1; r100; cheng2025empowering), have achieved significant advances in natural language processing (NLP)  (r10; r11; r12). However, with the advanced capabilities of LLMs, they are subject to frequent misused in various domains, including academic fraud, the creation of deceptive material, and the generation of fabricated information (r13; r14; r15; chen-etal-2025-llms-recognize), which underscores the critical need to distinguish between human-written text (HWT) and LLM-generated text (LGT) (r14; r18; r19).
However, due to their sophisticated functionality, LLMs pose significant challenges in the robustness of current AI detection systems (r12). The existing detection systems, including commercial ones, frequently misclassify texts as HWT (r20; r21) and generate inconsistent results when analyzing the same text using different detectors (r22; r23). Studies show false positive rates reaching up to 50% and false negative rates as high as 100% in different tools (r23) when dealing with out-of-distribution (OOD) datasets.
Another critical issue with the existing AI detection systems is their lack of verifiable evidence (r24; r200), as these tools typically provide only simple outputs like "likely written by AI" or percentage-based predictions (r23). The lack of evidence prevents users from defending themselves against false accusations (r22) and hinders organizations from making judgments based solely on the detection results without convincing evidences (r23). This problem is particularly troublesome not only because the low accuracy of such systems as mentioned before, but also due to the consequent inadequate response to LLM misuse, which can lead to significant societal harm (r25; r26; r27; r12). These limitations highlight the pressing need for more reliable, explainable and robust detectors.
@@
Empirical results show that IPAD outperforms state-of-the-art baselines by 9.05% in Average Recall on in-distribution datasets, 12.93% in AUROC on out-of-distribution (OOD) datasets, and 5.48% in AUROC under adversarial attacks. IPAD also generalizes well to structured data. A user study further reveals that IPAD improves trust and usability in detection tasks by presenting concrete decision evidence, including predicted prompts and regenerated texts. Code is available at https://github.com/Bellafc/IPAD-Inver-Prompt-for-AI-Detection.
@@ 2.1 Preliminaries
Our method comprises a Prompt Inverter  $f_{\text{inv}}$, and two Distinguishers, namely the Prompt-Text Consistency Verifier (PTCV) $f_{\text{PTCV}}$ and the Regeneration Comparator (RC) $f_{\text{RC}}$. Given an input text $T$, the task is to determine whether it is human-written (HWT) or generated by an LLM (LGT). We denote by $\mathcal{D}_{\text{PI}}$ the training set for $f_{\text{inv}}$, consisting of pairs $(T,P)$ where $T$ is an LLM-generated text and $P$ is its original prompt. The two distinguishers are trained using disjoint datasets: $\mathcal{D}_{\text{LGT}}$ contains LLM-generated samples and $\mathcal{D}_{\text{HWT}}$ contains human-written ones. All components are fine-tuned using Microsoft’s Phi3-medium-128k-instruct modelDBLP:journals/corr/abs-2404-14219.
@@
To estimate the fine-tuned model’s binary classification probability (i.e., the probability of predicting “yes” or “no”), we follow the logit-based estimation approach (yoshikawa-okazaki-2023-selective). Given the model input $x$, and the output logits $z$, the model’s probability assigned to $\hat{y}$ is computed through the softmax function $\sigma$ :
$\text{Confidence}_{\text{yes}}=P(\hat{y}=\text{``yes''}\mid x)=\sigma(z)_{\text{yes}},\quad\text{Confidence}_{\text{no}}=P(\hat{y}=\text{``no''}\mid x)=\sigma(z)_{\text{no}}$
@@
$\text{Probability}_{\text{yes}}=\frac{\text{Confidence}_{\text{yes}}}{\text{Confidence}_{\text{yes}}+\text{Confidence}_{\text{no}}},\quad\text{Probability}_{\text{no}}=1-\text{Probability}_{\text{yes}}$
@@ 2.2 Workflow
We perform inference on unseen input texts $T$ by sequentially applying the trained modules. Given an input text $T$, we first use the prompt inverter $f_{\text{inv}}$ to recover the most plausible prompt $P$. The prompt is then used to regenerate a candidate text $T^{\prime}$ via the an LLM. Next, we compute two probability scores: $p_{\text{PTCV}}$, indicating whether $T$ is consistent with $P$, and $p_{\text{RC}}$, assessing the likelihood that $T$ and $T^{\prime}$ originate from the same prompt. These scores are fused into a final decision score $\hat{p}$ using the gird-searched weight $w$, and the predicted label is determined by comparing $\hat{p}$ against the threshold $\tau$. The complete inference pipeline is summarized in Algorithm 1.
Algorithm 1 IPAD Detection Procedure
[…]
9:if $\hat{p}>\tau$ then
10:  $\hat{Y}\leftarrow\mathrm{LGT}$
11:else
12:  $\hat{Y}\leftarrow\mathrm{HWT}$
13:end if
[…]
2.3 Computational Complexity and Deployment Considerations
The inference procedure of the IPAD framework consists of three calls through a light-weight open-sourced LLM phi-3-medium-128k-instruct. Phi-3 is a decoder-only Transformer, whithin which, the self-attention complexity per layer is $\mathcal{O}(n^{2}\cdot d)$, where $n$ is the sequence length and $d$ is the hidden dimension (vaswani2017attention). The additional api call to gpt-3.5-turbo for re-generating texts introduces fixed latency but no local computation cost. Therefore, the overall computational cost is bounded by $\mathcal{O}(3\cdot L\cdot n^{2}\cdot d+\texttt{OpenAI}_{\text{api}})$, where $L=32$ is the number of layers in phi-3 (DBLP:journals/corr/abs-2404-14219), which is relatively small. All three phi-3 calls can be deployed in an Nvidia V100 GPU as the minimum requirement. This demonstrates that IPAD is not computationally expensive and can be deployed with relatively modest hardware requirements.
@@ 2.4 Training
The supervised fine-tuning  (r70) process is performed on a Microsoft’s open model, phi3-medium-128k-instruct, and we use low-rank adaptation (LoRA) method  (r69) on the LLaMA-Factory framework (r68). We train it using six A800 GPUs for 20 hours for Prompt Inverter, 7 hours for PTCV, and 9 hours for RC.
@@ 3.1.1 Evaluation Baselines and Metrics
The in-distribution experiments refer to the testing results presented in  (r3), where the data aligns with the training data used for the IPAD, thereby serving as our baseline. This baseline assesses how the RoBERTa classifiers (base and large) (park2021klue), the HC3 detector (hc3), and the OUTFOX detector (r3) perform on standard data as well as under DIPPER (alkanhel2023dipper) and OUTFOX attacks.
The OOD experiments refer to the DetectRL baseline  (r58), which is a comprehensive benchmark, which includes four datasets: (1) academic abstracts from the arXiv Archive (covering the years 2002 to 2017) , (2) news articles from the XSum dataset  (r59), (3) creative stories from Writing Prompts  (r60), and (4) social reviews from Yelp Reviews (r61). It also employs three attack methods to simulate complex real-world detection scenarios, which include (1) the prompt attacks, (2) paraphrase attacks, and (3) perturbation attacks (r58). DetectRL evaluates three classifiers on the OOD dataset: DetectLLM (LRR) (su2023detectllm), Fast-DetectGPT (bao2023fast), RoBERTa Classifier (Base). We included two more strong classifiers in our evaluation DetectLLM (NPR) (su2023detectllm) and Binoculars  (hans2024spotting). All the testing sets have 1,000 samples in our experiments.
We further evaluate its performance on OOD datasets with structured prompts. The LongWriter dataset (bai2025longwriter), featuring an average prompt length of 1,501 tokens, reflects IPAD’s capability to handle long-form prompts. The Code-Feedback(r101) and Math datasets (hendrycks2021measuring) contain highly structured prompts, in contrast to typical essay-like writing. We compare IPAD with baseline detectors from DetectRL to assess its relative performance under these challenging conditions.
The  Area Under Receiver Operating Characteristic curve (AUROC) is widely used for assessing detection method  (r55). Since our models predict binary labels, we follow the  Wilcoxon-Mann-Whitney statistic (r56), and the formula is shown in Appendix  B. The  AvgRec is the average of  HumanRec and  MachineRec, which refers to the recall of the Human-written texts and the LLM-generated texts (r57).
@@ 3.1.3 Comparison of IPAD with other detectors in and out of distribution
Figure 2: The In-distribution data performance of IPAD and the baseline detectors. Since  r3 only presents the AvgRec data for the baselines, we also calculate AvgRec data for IPAD to compare.
@@
Table 2: Detection Accuracy (AUROC %) on four diverse OOD datasets
@@
Table 3: Detection Accuracy (AUROC %) on three attacked OOD datasets
@@ 3.2.1 Necessity
To prove that it is necessary to fine-tune on IPAD with IPAD with  PTCV and  RC, we conducted ablation study to use the same finetune method on only  input texts and only  predicted prompts, with the finetune data format shown in Appendix  C. We only experimented on Prompt Inverter + PTCV and Prompt Inverter + RC to compare with the three-moduled IPAD.
@@ 3.2.2 Effectiveness
We use DPIC (r62) and PE (r65) as baseline methods for prompt extraction. DPIC employs a zero-shot approach using the prompt states in Appendix  D, while PE uses adversarial attacks to recover system prompts. In our evaluation, we tested 1000 LGT and 1000 HWT samples. We use only in-distribution data for testing since only these datasets include original prompts. The metrics are all tested on comparing the similarity of the original prompts and the predicted prompts. The results shown in Table  5 illustrate that IPAD consistently outperforms both DPIC and PE across all four metrics (BartScore (r64), Sentence-Bert Cosine Similarity (r63), BLEU (r66), and ROUGE-1 (r67)), which highlight the effectiveness of the IPAD  Prompt Inverter.
@@
We conducted a comparison study using the frozen Prompt Inverter but different distinguishing methods. The first and second methods employed Sentence-Bert (r63) and Bart-large-cnn (r64) to compute the similarity score between the input texts and the regenerated texts. We selected thresholds that maximized AvgRec, which were 0.67 for Sentence-Bert and -2.52 for Bart-large-cnn. The classification rule is that the texts with scores greater than the threshold will be classified as LGT, while the texts with scores less than or equal to the threshold will be classified as HWT. The third method is to directly prompt ChatGPT in Appendix  C, which mimic the fine-tuning process of PTCV and RC. The final results shown in Table  6 demonstrate that the other distinguishing methods performed worse than IPAD, highlighting the superior effectiveness of the IPAD  Distinguishers.
@@
DPIC first uses a zero-shot prompt inverter to generate prompts, then applies a Siamese encoder and classifier to measure similarity between the embeddings of the original and regenerated texts. However, the classifier’s reliance on embedding similarity is ambiguous, as similar texts may stem from different prompts. IPAD addresses this by fine-tuning directly on raw texts and reformulating the task as a logical reasoning problem as shown in the instructions of PTCV and RC. Our trained Prompt Inverter outperforms DPIC’s generic zero-shot method as shown in Table 5, and IPAD also achieves better performance than DPIC overall, as results shown in Appendix  E.